For spring break some students flew to cancun. From mexico city, the airplane took 2 hrs to fly 600 mi against a head wind. The return trip with the wind took 1 2/3 hr. Find the speed of the plane in still air.
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

600/2=300 MPH. AGAINST THE WIND.
600/(5/3)=600*3/5=1800/5=360 MPH WITH THE WIND.
(360-300)/3=60/2=30 MPH IS THE WIND SPEED.
360-30=330 THE PLANE'S SPEED IN STILL AIR.
300+30=330 DITTO.

Answer by Alan3354(69443)

The "plane's speed in still air" is called airspeed.